OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x2 = −4x − 5?

OpenStudy (amistre64):

an x intercept has y=0 therefore this equation is of the same nature as 0 - x^2 = -4x -5 0 = x^2 -4x -5 which can be solved with a few different methods

OpenStudy (anonymous):

Thanks ! I was stuck.

OpenStudy (amistre64):

were you stuck at getting it into that form? or in solving it from that form?

OpenStudy (amistre64):

the simplest approach to the solution of a quadratic equation is known as the quadratic formula:\[x=\frac{-b}{2a}\pm \frac{\sqrt{b^2-4ac}}{2a}\] given the equation in the form: ax^2 + bx + c

OpenStudy (anonymous):

I was stuck solving it. But I got it, the answer was -1,0 and 5,0.

OpenStudy (amistre64):

good job then ;)

OpenStudy (anonymous):

Thanks